A trapezoid with an area of 48m has a height of 6m. One of the bases is 12m. I need to solve the equation to find the length of its other base?
Answer by rothauserc(4718)

A = (1/2)h(B+b) where A is area of the trapezoid, h is height, B is one base, b is the other base, therefore
48 = (1/2)6(12+b)
48 = 3(12+b)
48 = 36 + 3b
3b = 12
b = 4
the other base is 4m